A club with 32 members is to select five officers. In how many ways can this be done?
Answer by stanbon(75887)

A club with 32 members is to select five officers. In how many ways can this be done?
--------
Order is not implied.
Ans: 32C5 = 32!/[(32-5)!*5!] = (32*31*30*29*28)/(1*2*3*4*5) = 201376
-----------------
Cheers,
Stan H.
====================